### v1

@@ Appendix A Supporting Lemmas
(54) $\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta\left(k+1\right)}\right\|^{2}\right]}$
$\leqslant\frac{1+\bar{\rho}_{W}}{2}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]}+\frac{8n\gamma^{2}\beta^{2}\bar{\rho}_{W}K}{1-\bar{\rho}_{W}}\sigma^{2}+\frac{4\gamma^{2}\beta^{2}\bar{\rho}_{W}}{1-\bar{\rho}_{W}}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]}.$
Adding $\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{0}\right\|^{2}\right]$ on both sides and noticing $\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]}\leqslant\sum_{k=0}^{K}{\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]}$, we can obtain the target result, which completes the proof.
Lemma 8 (Accumulated tracking error).
Suppose Assumptions 24 hold. Let the stepsize satisfy (49). Then, for all $K\geqslant 1$,
(55) $\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]}$
$\leqslant\frac{4\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{0}\right\|^{2}\right]}{\left(1-\bar{\rho}_{W}\right)K}+\frac{24n\sigma^{2}}{\left(1-\bar{\rho}_{W}\right)K}+\frac{96\bar{\rho}_{W}L^{2}}{\left(1-\bar{\rho}_{W}\right)^{2}}\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]}$
@@
(56) $\sum_{k=0}^{K}{\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]}$ $\leqslant\frac{3+\bar{\rho}_{W}}{4}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]}+\frac{18\bar{\rho}_{W}L^{2}}{1-\bar{\rho}_{W}}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]}$
@@
we can obtain the target result.

### v2

@@ 6 Conclusions
$\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta\left(k+1\right)}\right\|^{2}\right]}$ (60)
$\leqslant\frac{1+\bar{\rho}_{W}}{2}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]}+\frac{8n\gamma^{2}\beta^{2}\bar{\rho}_{W}K}{1-\bar{\rho}_{W}}\sigma^{2}$
$\quad+\frac{4\gamma^{2}\beta^{2}\bar{\rho}_{W}}{1-\bar{\rho}_{W}}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]}.$
Adding $\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{0}\right\|^{2}\right]$ on both sides and noticing that
$\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]}\leqslant\sum_{k=0}^{K}{\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]},$
we can obtain the target result, which completes the proof.
Lemma 8 (Accumulated GT error)
Suppose Assumptions 24 hold. Let the stepsize satisfy (55). Then, for all $K\geqslant 1$, we have
$\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]}$ (61)
$\leqslant\frac{4\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{0}\right\|^{2}\right]}{\left(1-\bar{\rho}_{W}\right)K}+\frac{24n\sigma^{2}}{\left(1-\bar{\rho}_{W}\right)K}$
$\quad+\frac{96\bar{\rho}_{W}L^{2}}{\left(1-\bar{\rho}_{W}\right)^{2}}\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]}$
@@
$\sum_{k=0}^{K}{\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]}$ (62)
$\leqslant\frac{3+\bar{\rho}_{W}}{4}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]}+\frac{18\bar{\rho}_{W}L^{2}}{1-\bar{\rho}_{W}}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]}$
@@
we can obtain the desired result.